OpenStudy (anonymous):

Simplify. √5 (8+3 √6)

Vocaloid (vocaloid):

distribute A(B+C) = A*B + A*C

OpenStudy (anonymous):

k

Vocaloid (vocaloid):

A(B+C) = A*B + A*C √5 (8+3 √6) = ____ ?